a one-way road passes under an overpass in the shape of half and ellipse, 30ft high at the center and 20ft wide. Assuming a truck is 12ft wide, what is the tallest truck that can pass under the overpass?
Found 2 solutions by lwsshak3, ikleyn:
Answer by lwsshak3(11628)

a one-way road passes under an overpass in the shape of half and ellipse, 30ft high at the center and 20ft wide. Assuming a truck is 12ft wide, what is the tallest truck that can pass under the overpass?
***
Ellipse has a vertical major axis
Its basic form of equation:
x^2/b^2+y^2/a^2=1
For given problem:
a=30
b=10
equation: x^2/10^2+y^2/30^2=1
=x^2/100+y^2/900=1
coordinates of point at which a 12 ft wide truck would reach the overpass=(6,y)
plug in x-coordinate and solve for y
y^2/900=1-x^2/100=1-36/100=54/100
y^2=9*54=486
y=√486≈22
what is the tallest truck that can pass? < 22 ft

Answer by ikleyn(53965)

.
a one-way road passes under an overpass in the shape of half and ellipse, 30ft high at the center and 20ft wide.
Assuming a truck is 12ft wide, what is the tallest truck that can pass under the overpass?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        Calculations in the post by @lwsshak3 are incorrect, leading to wrong answer.
        I came to provide a correct solution.


Ellipse has a vertical major axis
Its basic form of equation:
x^2/b^2+y^2/a^2=1
For given problem:
a=30
b=10
equation: x^2/10^2+y^2/30^2=1
=x^2/100+y^2/900=1
coordinates of point at which a 12 ft wide truck would reach the overpass=(6,y)
plug in x-coordinate and solve for y
y^2/900=1-x^2/100=1-36/100=64/100         <<<---=== this is the point where @lwsshak3 made his mistake
y^2=9*64=576
y=√576 = 24

what is the tallest truck that can pass? < 24 ft


More accurately, the conception "tallest truck" does not work in this problem,
so the question is posed incorrectly in the post.

A correct question should ask about the set of possible heights.
Then the answer " h < 24 ft " is correct and accurate.


Solved correctly.